# TetSphere Splatting: Representing High-Quality Geometry with Lagrangian Volumetric Meshes

### v2

@@ 4 TetSphere Splatting
We employ tetrahedral spheres as our primitive of choice. Unlike point clouds, tetrahedral meshes enforce structured local connectivity between points owing to tetrahedralization. This preserves the geometric integrity of the 3D shape and also enhances the surface quality by imposing geometric regularization across the entire mesh interior. We formulate the reconstruction of shapes through TetSphere splatting as a deformation of tetrahedron spheres. Starting from a set of tetrahedral spheres, we adjust the positions of their vertices to align the rendered images of these meshes with the corresponding target multi-view images. Vertex movement is constrained by two geometric regularizations on the tetrahedral meshes, derived from the field of geometry processing [4]. These constraints, which penalize the nonsmooth deformation (via bi-harmonic energy) and prevent the inversion of mesh elements (via local injectivity), have proven effective in ensuring that the resulting tetrahedral meshes are of superior quality and maintain structural integrity. Fig. 4 illustrates the iterative process of TetSphere splatting with deforming tetrahedral spheres.
4.1 Tetrahedral Sphere Primitive
The primitive of TetSphere splatting is a tetrahedralized sphere, called TetSphere, with $N$ vertices and $T$ tetrahedra. By applying principles from the Finite Element Method (FEM) [79], the mesh of each sphere is composed of tetrahedral elements, with each tetrahedron constituting a 3D discrete piecewise linear volumetric entity. We denote the position vector of all vertices of the $i$-th deformed sphere mesh as ${x}_{i}\in\mathbb{R}^{3N}$. The deformation gradient of the $j$-th tetrahedron in the $i$-th sphere is denoted as $\mathbf{F}^{(i,j)}_{\mathbf{x}}\in\mathbb{R}^{3\times 3}$, which quantitatively describes how each tetrahedron’s shape transforms [79]. Essentially, the deformation gradient $\mathbf{F}^{(i,j)}_{\mathbf{x}}$ serves as a measure of the spatial changes a tetrahedron undergoes from its original configuration to its deformed state. Refer to the inset figure for a visual explanation and Appendix 0.D for an in-depth derivation.
Rather than using a single sphere, our representation employs a collection of spheres to accurately represent arbitrary shapes. Consequently, the complete shape is the union of all spheres. By adopting multiple spheres, this approach ensures that each local region of a shape is detailed independently, enabling a highly accurate representation. Moreover, it allows for the representation of shapes with arbitrary topologies. Such a claim is theoretically guaranteed by the paracompactness property of manifold shapes [37].
  • Compared to neural representations (e.g., NeRF, NeuS), our tetrahedral representation does not rely on neural networks, thus inherently accelerating the optimization process.
  • Compared to Eulerian representations (such as DMTet), our approach entirely avoids the need for iso-surface extraction – an operation that often degrades mesh quality owing to the predetermined resolution of the grid space. Furthermore, Eulerian methods rely on level-set functions, a dependency that can lead to undesirable floating and noisy artifacts given sparsely populated or unpopulated grid values during the training phase. The deformation process of our representation inherently prevents individual spheres from breaking, thereby eliminating floating noises.
  • Compared to other Lagrangian representations, such as triangle meshes and Gaussian point clouds, our method offers a volumetric representation through the use of tetrahedral meshes. It is more robust to thin shapes, which often pose a risk of self-penetration in surface meshes, and to fragile structures, which are a challenge for point clouds. Furthermore, unlike Gaussian point clouds, which require additional steps to produce a mesh, our tetrahedral representation naturally forms a mesh. Each tetrahedron also imposes constraints among vertices, leading to superior mesh quality.
4.2 TetSphere Splatting as Shape Deformation
To reconstruct the geometry of the target object, we deform the initial TetSpheres by changing their vertex positions. This process is governed by two primary goals: ensuring the deformed TetSpheres align with the input multi-view images and maintaining high mesh quality that adheres to necessary geometry constraints.
To maintain the mesh quality, we leverage bi-harmonic energy – defined in the literature on geometry processing [6] as an energy quantifying smoothness throughout a field – to the deformation gradient field. This geometric regularization ensures the smoothness of the deformation gradient field across the deformation process, thus preventing irregular mesh or bumpy surfaces. It is important to highlight that this bi-harmonic regularization does not lead to over-smoothness of the final result. This is because the energy targets the deformation gradient field, which measures the relative changes in vertex positions, rather than the absolute positions themselves. Such an approach allows for the preservation of sharp local geometric details, akin to techniques used in physical simulations [95]. Furthermore, we introduce a geometric constraint to guarantee local injectivity in all deformed elements [74]. This ensures that the elements maintain their orientation during the deformation, avoiding inversions or inside-out configurations. This constraint can be mathematically expressed as $\mathrm{det}(\mathbf{F}^{(i,j)}_{\mathbf{x}})>0.$ Importantly, these two terms – bi-harmonic energy for smoothness and local injectivity for element orientation – are universally applicable to any tetrahedral meshes, stemming from their fundamental basis in geometry processing [4].
@@ 4.2 TetSphere Splatting as Shape Deformation
Let $\mathbf{x}=[x_{1},...,x_{M}]\in\mathbb{R}^{3NM}$ denote the positions of vertices across all $M$ TetSpheres, and $\mathbf{F}_{\mathbf{x}}\in\mathbb{R}^{9MT}=[\mathrm{vec}(\mathbf{F}^{(1,1)}_{\mathbf{x}}),...,\mathrm{vec}(\mathbf{F}^{(M,T)}_{\mathbf{x}})]$ denote the flattened deformation gradient fields of all TetSpheres. In the bi-harmonic energy, the Laplacian matrix is defined based on the connectivity of the tetrahedron faces, denoted as $\mathbf{L}\in\mathbb{R}^{9MT\times 9MT}$. This matrix is block symmetric, where each block $\mathbf{L}_{pq}\in\mathbb{R}^{9\times 9},p\neq q$ is set to a negative identity matrix $-I$ if the $p$-th and $q$-th tetrahedron shares a common triangle; or $kI$ for $\mathbf{L}_{pp}$, where $k$ is the number of neighbors of the $p$-th tetrahedron. The deformation of the TetSpheres is formulated as an optimization problem:
@@
In the proposed optimization framework, three considerations have been outlined. 1) The adaptive loss function $\mathbf{\Phi}(\cdot)$, designed to be flexible, supports a variety of metrics, including $l_{1}$ for color images, MSE for depth images, cosine embedding loss for normal images, or SDS loss. 2) Given that the tetrahedron is a linear element, the deformation gradient $\mathbf{F}^{(i,j)}_{\mathbf{x}}$ is a linear function of $\mathbf{x}$, making the bi-harmonic energy a quadratic term. 3) The weights $w_{1}$ and $w_{2}$ are dynamically adjusted using a cosine scheduler. We provide details of the scheduler’s hyperparameters in Appendix 0.A.
5 Optimization Framework
The overall framework of optimization is illustrated in Fig. 3. The geometry optimization follows the shape deformation as detailed in Sec. 4. In this section, we introduce the remaining two components: initialization of TetSphere splatting and texture optimization.
5.1 TetSphere Initialization
@@
Given multi-view images as inputs, we select feature points to initialize the 3D center positions of the TetSpheres. We aim to achieve a uniform distribution of these TetSpheres within the object, ensuring comprehensive coverage of the silhouette depicted in the multi-view images.
Figure 5: Feature point selection via silhouette coverage algorithm. Our method automates the feature point selection from a sparse voxel constructed from multi-view images. By solving a linear programming problem, we determine these feature points and initialize tetrahedral spheres for subsequent geometry optimization.
@@
We introduce an algorithm, silhouette coverage, inspired by Coverage Axis [18] to automatically select initial centers of TetSpheres for an arbitrary shape. This process begins with constructing a coarse voxel grid and initially assigning a zero value to each voxel. By projecting these voxels into the image spaces using the same camera poses as the input multi-view images, voxels within the foreground of all images are marked with a value of $1$. These voxel positions are identified as candidate positions of TetSphere centers. From these marked positions, the objective is to pick a minimal subset of candidates that ensure thatall candidates are fully encapsulated by TetSpheres centered on these points. This involves placing uniform spheres with varying radius values at all candidate points and choosing a minimal subset that collectively covers all the candidate points. We formulate a linear programming problem to efficiently perform the selection. The detailed formulation is provided in Appendix 0.E. Fig. 5 shows the pipeline of the silhouette coverage algorithm. In our implementation, with a voxel grid resolution $300\times 300$ and $n=20$, the whole TetSphere initialization completes in ${\sim}1$ minute on average.
5.2 Texture/PBR Material Optimization
The final stage of our reconstruction process focuses on optimizing the textures or material properties. Our method, with its explicit representation, allows textures and materials to be directly applied to the surface vertices and faces of the TetSpheres. This enables the use of sophisticated material models, such as Disney’s principled BRDF [7], with physically-based rendering.
Material optimization is facilitated through the use of differentiable rasterizers [43], which adjust the textures or materials to closely match the input multi-view color images. A significant advantage of TetSphere splatting is that the deformation of tetrahedral spheres does not alter the surface topology. Unlike methods such as DMTet, which require isosurface extraction and subsequent texture parameterization at each step due to potential changes in the underlying shape, our method necessitates only a single texture parameterization at the beginning of optimization. This parameterization remains consistent throughout the process, significantly enhancing the efficiency of texture optimization.
For scenarios with dense input views, we have found that using textured images as optimization variables is straightforward and yields high-quality results. In cases with sparse input views, we adopt a two-layer multilayer perceptron (MLP) that takes the surface vertex positions as inputs and outputs the material parameters, a practice in line with existing methods [80, 69].
Table 1: Single-View reconstruction results on the GSO Dataset: Evaluating reconstruction accuracy with Chamfer Distance (Cham.) and Volume IoU, alongside new mesh quality metrics: Area-Length Ratio (ALR), Manifoldness Rate (MR), and Connected Component Discrepancy (CC Diff.). TetSphere splatting demonstrates superior performance on these criteria and maintains competitive reconstruction accuracy.
[…]
Realfusion [56] Marching Cubes 0.0819 0.2741 0.0561 100% 47.7
Magic123 [68] Marching Tets 0.0516 0.4528 0.0383 100% 13.7
One-2-3-45 [47] Marching Cubes 0.0629 0.4086 0.0574 96% 0.83
Point-E [64] Marching Cubes 0.0426 0.2875 0.2421 100% 18.38
Shap-E [39] Marching Tets 0.0436 0.3584 0.1236 100% 9.03
Zero123 [49] Marching Cubes 0.0339 0.5035 0.0543 100% 0.18
SyncDreamer [51] Marching Cubes 0.0261 0.5421 0.0201 10% 0.3
Wonder3d [53] Marching Cubes 0.0329 0.5768 0.0281 100% 0.0
Open-LRM [31] Marching Cubes 0.0285 0.5945 0.0252 100% 0.0
DreamGaussian [83] Marching Cubes 0.0641 0.3476 0.0812 100% 237.4
Ours N/A 0.0351 0.6317 0.3665 100% 0.0
Figure 6: Qualitative comparison on image-to-3D generation with surface mesh visualizations and rendered normal maps (1/2). Our generated meshes exhibit minimal noise and high quality, featuring more regular triangle meshing.
Figure 7: Qualitative comparison on image-to-3D generation with surface mesh visualizations and rendered normal maps (2/2). Our generated meshes exhibit minimal noise and high quality, featuring more regular triangle meshing.
@@ 6 Experiments and Results
We conduct experiments on three applications to demonstrate TetSphere’s versatility and effectiveness: single-view reconstruction, image-to-3D shape generation, and text-to-3D shape generation. We choose single-view reconstruction because it allows for comprehensive evaluation using the well-established benchmark on the Google Scanned Objects (GSO) dataset. We choose the latter two applications because the existing methods are known for their high computational costs, specifically in terms of GPU quantity and memory capacity. We conduct experiments on these applications to demonstrate that TetSphere splatting effectively addresses and mitigates these issues.
For the first two applications, the input is a single image from which a textured 3D shape is reconstructed. For text-to-3D generation, the input is a text prompt. All the output targets are textured 3D shapes. For single-view reconstruction, we conduct a quantitative evaluation of our method on GSO dataset, comparing its performance with other state-of-the-art methods. The evaluation focuses on both the reconstruction accuracy and the geometry quality. In addition, we show a series of qualitative comparisons using web-collected images or text prompts. Appendix 0.A provides the implementation details.
@@ 6.1 Baselines and Evaluation Protocol
6.1.1 Baselines.
For single-view reconstruction, we quantitatively compare with several state-of-the-art methods: RealFusion [56], Magic123 [68], One-2-3-45 [47], Point-E [64], Shap-E [39], Zero123 [49], SyncDreamer [50], Wonder3d [53], Open-LRM [33, 31], and DreamGaussian [83]. For image-to-3D generation, we compare TetSphere splatting with Magic123 and Dreamcraft3D [80]. Both are multi-stage methods starting with NeRF optimization followed by DMTet for optimizing mesh and texture. For text-to-3D, our comparison focuses on RichDreamer, a state-of-the-art method known for incorporating PBR materials and generating shapes, showcasing notable results in 3D generation from text prompts.
6.1.2 Evaluation Datasets.
For single-view reconstruction, following prior research [51, 53], we use the GSO dataset for our evaluation, which covers a broad range of everyday objects. The evaluation dataset aligns with those used by SyncDreamer and Wonder3D, featuring 30 diverse objects ranging from common household items to animals. For image-to-3D generation, we include a variety of internet-collected images in different styles in our evaluation. For text-to-3D generation, we employ text prompts from the official implementation of RichDreamer.
6.1.3 Metrics.
@@
To assess the accuracy of single-view reconstruction, we use two commonly used metrics: Chamfer Distance (Cham.) and Volume IoU, comparing the ground-truth shapes with the reconstructed ones. Following established practices, we use the rigid Iterative Closest Point (ICP) [3] algorithm to align the generated shapes with their ground-truth counterparts before metric calculation.
@@
While Cham. and Volume IoU effectively gauge volumetric and point-based shape conformity, they fall short of evaluating mesh quality. Recognizing this gap, we introduce three additional metrics to comprehensively assess the geometry quality of the generated shapes: 1) Area-length Ratio (ALR): This metric computes the average ratio of a triangle’s area to its perimeter (scaled by a constant coefficient) within the surface mesh. Values range from 00 to $1$, where meshes with higher ALR values contain mostly equilateral triangles, thereby indicating superior triangle quality; 2) Manifoldness Rate (MR): Manifoldness verifies whether a mesh qualifies as a closed manifold. Non-manifold meshes can manifest anomalies, such as edges shared by more than two faces, vertices connected by edges but not by a surface, isolated vertices and edges, or open boundaries, which can cause significant problems in downstream applications such as simulation and rendering. We report the percentage of manifold shapes within the evaluation dataset as MR; and 3) Connected Component Discrepancy (CC Diff.) from the ground-truth shape: This measure identifies the presence of floating artifacts or structural discontinuities within the mesh, highlighting the integrity and cohesion of the reconstructed shape.
@@
NeRF Make-it-3D[84] 4 1.22
Magic123[68] 4 1.03
NeuS SyncDreamer[50] 48 1.8
DMTet DreamCraft3D[80] 8 1.23
GeoDream[54] 8 1.30
3D GS DreamGaussian[83] 80 4.43
Ours 120 6.59
Figure 8: Analysis on energy coefficients for geometry optimization. Dark regions indicate flipping of the surface triangles.

### v4

@@ 3.2 TetSphere Splatting as Shape Deformation
Let $\mathbf{x}=[x_{1},...,x_{M}]\in\mathbb{R}^{3NM}$ denote the positions of vertices across all $M$ TetSpheres, and $\mathbf{F}_{\mathbf{x}}\in\mathbb{R}^{9MT}=[\mathrm{vec}(\mathbf{F}^{(1,1)}_{\mathbf{x}}),...,\mathrm{vec}(\mathbf{F}^{(M,T)}_{\mathbf{x}})]$ denote the flattened deformation gradient fields of all TetSpheres, where $N$ and $T$ denote the number of vertices and tetrahedra within each TetSphere, respectively. In the bi-harmonic energy, the Laplacian matrix is defined based on the connectivity of the tetrahedron faces, denoted as $\mathbf{L}\in\mathbb{R}^{9MT\times 9MT}$. This matrix is block symmetric, where each block $\mathbf{L}_{pq}\in\mathbb{R}^{9\times 9},p\neq q$ is set to a negative identity matrix $-I$ if the $p$-th and $q$-th tetrahedron shares a common triangle; or $kI$ for $\mathbf{L}_{pp}$, where $k$ is the number of neighbors of the $p$-th tetrahedron. The deformation of the TetSpheres is formulated as an optimization problem:
@@
In the proposed optimization framework, three considerations have been outlined. 1) The adaptive loss function $\mathbf{\Phi}(\cdot)$, designed to be flexible, supports a variety of metrics, including $l_{1}$ for color images, MSE for depth images, and cosine embedding loss for normal images. These losses can all be optimized using gradients provided by differentiable renderers, such as mesh rasterizers (Laine et al., 2020).. 2) Given that the tetrahedron is a linear element, the deformation gradient $\mathbf{F}^{(i,j)}_{\mathbf{x}}$ is a linear function of $\mathbf{x}$, making the bi-harmonic energy a quadratic term. 3) The weights $w_{1}$ and $w_{2}$ are dynamically adjusted using a cosine scheduler. We provide details of the scheduler’s hyperparameters in Appendix I.
4 TetSphere Initialization and Texture Optimization
@@
TetSphere initialization. Given multi-view images as inputs, we select feature points to initialize the 3D center positions of the TetSpheres. We aim to achieve a uniform distribution of these TetSpheres, ensuring comprehensive coverage of the silhouette depicted in the multi-view images.
@@
We introduce an algorithm, silhouette coverage, inspired by Coverage Axis (Dou et al., 2022) to select initial centers of TetSpheres for an arbitrary shape automatically. This process begins with constructing a coarse voxel grid and initially assigning a zero value to each voxel. By projecting these voxels into the image spaces using the same camera poses as the input multi-view images, voxels within the foreground of all images are marked with a value of $1$. These voxel positions are identified as candidate positions of TetSphere centers. From these marked positions, the objective is to pick a minimal subset of candidates to ensure that all candidates are fully encapsulated by TetSpheres and centered on these points. This involves placing uniform spheres with varying radius values at all candidate points and choosing a minimal subset that collectively covers all the candidate points. We formulate a linear programming problem to perform the selection efficiently. The detailed formulation is provided in Appendix G. In our implementation, with a voxel grid resolution $300\times 300$ and $M=20$, the whole TetSphere initialization completes in ${\sim}1$ minute on average.
Texture optimization. While the primary purpose of TetSphere splatting is to represent high-quality geometry, its explicit structure allows textures and materials to be directly applied to the surface vertices and faces of the TetSpheres. A key advantage of TetSphere splatting is that the deformation of tetrahedral spheres preserves the surface topology, enabling the seamless integration of advanced material models, such as Disney’s principled BRDF (Burley & Studios, 2012), in physically-based rendering for applications like text-to-3D generation. However, it is important to note that texture optimization is an optional feature of our representation. The primary objective of TetSphere splatting remains the accurate representation of geometry, with textures and materials being secondary enhancements. Further details are provided in Appendix H.
@@ 5 Experiments and Results
We provide detailed quantitative evaluations on 3D reconstruction from both multi-view and single-view images to demonstrate the effectiveness of TetSphere. Additionally, we showcase TetSphere’s versatility through case studies on image-to-3D and text-to-3D shape generation. We further analyze the effect of energy coefficients and offer a comparative study on computational costs, particularly focusing on GPU requirements and memory usage for image-to-3D generation with SDS loss, where existing methods struggle with high computational demands.
5.1 Implementation Details
Our implementation of the TetSphere initialization algorithm is developed in C++ and uses the Gurobi linear programming solver. The optimization of geometric energies is implemented using CUDA as a PyTorch extension to enhance computational efficiency. 1) For multi-view reconstruction, we render multi-view RGBA and depth images of the ground truth shape as the input of TetSphere splatting. The optimization objective $\mathbf{\Phi}(\cdot)$ includes both rendering loss (MSE on the alpha mask) and depth loss, following Son et al. (2024). 2) For single-view reconstruction, we use Wonder3d (Long et al., 2023) to generate six multi-view images with predefined camera poses. The optimization objective $\mathbf{\Phi}(\cdot)$ includes rendering loss $l_{1}$ norm on tone-mapped color and MSE on the alpha mask) and normal loss (cosine loss on normals), following Munkberg et al. (2022). The implementation details for the two applications are described in Appendix I. To enhance robustness and efficiency, we apply a cosine scheduler to scale the coefficients of the geometry loss, formulated as $\eta=4^{\sin(\frac{t\pi}{2n})}$, where $t$ denotes the current iteration and $T$ represents the total number of iterations.
Table 1: Multi-View reconstruction results: Evaluating reconstruction accuracy with Chamfer Distance (Cham.) and Volume IoU, alongside mesh quality metrics: Area-Length Ratio (ALR), Manifoldness Rate (MR), and Connected Component Discrepancy (CC Diff.). For additional results on other metrics, please refer to Table 5 in Appendix E.1.
Method Geo. Rep. Cham. $\downarrow$ Vol. IoU $\uparrow$ ALR $\uparrow$ MR(%) $\uparrow$ CC Diff. $\downarrow$
NIE Eulerian 0.0254 0.1863 0.0273 72.3 7.0
FlexiCubes Eulerian 0.0247 0.5887 0.0722 45.5 201.3
NeuS Eulerian 0.0192 0.6182 0.0573 72.3 8.1
VolSDF Eulerian 0.0185 0.6423 0.0622 81.8 7.3
2DGS Lagrangian 0.0322 0.4923 0.0209 27.3 25.1
DMesh Lagrangian 0.0136 0.5616 0.1193 9.09 3.75
Ours Lagrangian 0.0184 0.6844 0.6602 100 0.0
@@ 5.2 Baselines and Evaluation Protocol
Baselines. For multi-view reconstruction, we quantitatively compare our method with state-of-the-art geometry representations: Neural Implicit Evolution (NIE) (Mehta et al., 2022), FlexiCubes (Shen et al., 2023b), 2DGS (Huang et al., 2024), and DMesh (Son et al., 2024). The first two are Eulerian representations, whereas the latter two are Lagrangian representations. For single-view reconstruction, we quantitatively compare with several state-of-the-art methods: Magic123 (Qian et al., 2023), One-2-3-45 (Liu et al., 2023a), SyncDreamer (Liu et al., 2023c), Wonder3d (Long et al., 2023), Open-LRM (Hong et al., 2023b; He & Wang, 2023), and DreamGaussian (Tang et al., 2023a). DreamGaussian is the only Lagrangian representation, as the field of single-view reconstruction is largely dominated by Eulerian methods. For all baselines, we follow their original implementations and use their publicly available codebase to get the results in this paper.
Evaluation datasets. For multi-view reconstruction, we follow Son et al. (2024) and use four closed-surface models from the Thingi32 dataset (Zhou & Jacobson, 2016), four open-surface models from the DeepFashion3D dataset (Heming et al., 2020), and two additional models with both closed and open surfaces from the Objaverse dataset (Deitke et al., 2023). We also add the challenging sorter shape from Google Scanned Objects (GSO) dataset (Downs et al., 2022) as it features slender structures. For single-view reconstruction, following prior research (Liu et al., 2023c; Long et al., 2023), we use the GSO dataset for our evaluation, which covers a broad range of everyday objects. The evaluation dataset aligns with those used by SyncDreamer and Wonder3D, featuring 30 diverse objects ranging from household items to animals.
Figure 5: Qualitative results on multi-view reconstruction, with surface mesh visualizations and rendered normal maps. Our method excels over baseline methods regarding mesh quality, less bumpy surface, correct surface orientation, and accurately capturing slender and thin structures.
Table 2: Single-View reconstruction results on the GSO Dataset: Evaluating reconstruction accuracy with Chamfer Distance (Cham.) and Volume IoU, alongside mesh quality metrics: Area-Length Ratio (ALR), Manifoldness Rate (MR), and Connected Component Discrepancy (CC Diff.).
Method Geo. Rep. Cham. $\downarrow$ Vol. IoU $\uparrow$ ALR $\uparrow$ MR(%) $\uparrow$ CC Diff. $\downarrow$
Magic123 Eulerian 0.0516 0.4528 0.0383 100 13.7
One-2-3-45 Eulerian 0.0629 0.4086 0.0574 96 0.83
SyncDreamer Eulerian 0.0261 0.5421 0.0201 10 0.3
Wonder3d Eulerian 0.0329 0.5768 0.0281 100 0.0
Open-LRM Eulerian 0.0285 0.5945 0.0252 100 0.0
DreamGaussian Lagrangian 0.0641 0.3476 0.0812 100 237.4
Ours Lagrangian 0.0351 0.6317 0.3665 100 0.0
@@
Metrics. To assess the accuracy of the reconstruction, we use two commonly used metrics: Chamfer Distance (Cham.) and Volume IoU (Vol. IoU), comparing the ground-truth shapes with the reconstructed ones. Following established practices, we use the rigid Iterative Closest Point (ICP) algorithm to align the generated shapes with their ground-truth counterparts before metric calculation. For multi-view reconstruction, we also report F-Score, Normal Consistency, Edge Chamfer Distance, and Edge F-Score following prior research (Shen et al., 2023b; Son et al., 2024).
@@
While these metrics effectively measure the reconstruction accuracy, they fail to evaluate mesh quality. We use three additional metrics to assess the geometry quality of the reconstructed shapes: 1) Area-length Ratio (ALR): This metric computes the average ratio of a triangle’s area to its perimeter (scaled by a constant coefficient) within the surface mesh. Values range from 00 to $1$, where meshes with higher ALR values contain mostly equilateral triangles, thereby indicating superior triangle quality; 2) Manifoldness Rate (MR): Manifoldness verifies whether a mesh qualifies as a closed manifold. Non-manifold meshes can manifest anomalies, such as edges shared by more than two faces, vertices connected by edges but not by a surface, isolated vertices and edges, and vertices where more than one distinct surface meets, which can cause problems in downstream applications such as simulation and rendering. We report the percentage of manifold shapes within the evaluation dataset as MR; and 3) Connected Component Discrepancy (CC Diff.) from the ground-truth shape: This measure identifies the presence of floating artifacts or structural discontinuities within the mesh, highlighting the integrity and cohesion of the reconstructed shape.
@@
Eulerian Make-it-3D 4 1.22
Magic123 4 1.03
SyncDreamer 48 1.8
DreamCraft3D 8 1.23
GeoDream 8 1.30
Lagrangian DreamGaussian 80 4.43
Lagrangian Ours 120 6.59
Figure 6: Analysis on geometry energy coefficients. Dark regions indicate flipping of the surface triangles.